OpenStudy (anonymous):

show that f(x)=.25*in(x) and g(x)=e^(4x) are inverse functions.

OpenStudy (anonymous):

Hint: if they are to be inverse, they must satisfy the property \(f(g(x))=g(f(x))=x\). Show that that equality holds for all \(x\).

OpenStudy (kropot72):

\[Let\ f(x) = y\] \[y=0.25\times \ln x\] To find f(x)^-1 interchange x and y and solve for y. \[x=0.25\times \ln y\] \[\ln y=4x ........(1)\] Can you solve e

OpenStudy (anonymous):

@kropot72, technically, that is not a proof, but rather a derivation, which is contrary to what the question is asking.

OpenStudy (kropot72):

Can you solve equation (1) to find y?

OpenStudy (anonymous):

it satisfys both methods. I need a proof, so ill use the first one. thanks to both of you.

OpenStudy (kropot72):

@yakeyglee The question does not ask for a proof but states "Show that.........". I had intended to point out that the inverse of the second expression should also be found to show that each expression was the inverse of the other.

OpenStudy (anonymous):

@kropot72 Questions that ask "Show that..." request a proof. Think about how you would approach these two questions differently: -Show that \(x=4\) is a solution to \(x^2-3x-4=0\). -Find a positive solution of \(x^2-3x-4=0\). The first requests a proof, the second requests a derivation. The first one is analogous to this problem.

OpenStudy (kropot72):

@yakeyglee You have made a good point :)